OpenStudy (anonymous):

Use the definition of laplace transform to find the transform of f (t)=4^t

zepdrix (zepdrix):

\[\Large\rm \mathscr{L}\left(4^t\right)=\int\limits^0^{\infty} 4^t e^{-st}dt\]Hey there :) So this is how we set up our definition, yes?

zepdrix (zepdrix):

Woah my boundaries tweeked out :O sec...

zepdrix (zepdrix):

\[\Large\rm \mathscr{L}\left(4^t\right)=\int\limits\limits_0^{\infty} 4^t e^{-st}dt\]

zepdrix (zepdrix):

If you change the base on the exponential, it will be easier to integrate. Recall that:\[\Large\rm \color{orangered}{a}=e^{\ln(\color{orangered}{a})}\]We want to apply this idea to the 4^t.

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{4^t}=e^{\ln(\color{orangered}{4^t})}\]

zepdrix (zepdrix):

And then applying your exponent rule:\[\Large\rm 4^t=e^{t \ln(4)}\]

zepdrix (zepdrix):

\[\Large\rm \mathscr{L}\left(4^t\right)=\int\limits\limits\limits_0^{\infty} e^{t \ln(4)}e^{-st}dt\]Again use rules of exponents to combine the terms :)

zepdrix (zepdrix):

\[\Large\rm \mathscr{L}\left(4^t\right)=\int\limits\limits\limits\limits_0^{\infty} e^{(\ln4-s)t}dt\]

zepdrix (zepdrix):

And it should be easier to work with from here I think.. hmm

ganeshie8 (ganeshie8):

this looks like a trick question as 4^t > e^t

zepdrix (zepdrix):

I think it's still ok. It still converges if we assume that \(\Large\rm s\gt \ln4\). Just need to give a proper restriction to s, yah? :o

ganeshie8 (ganeshie8):

Exactly !

zepdrix (zepdrix):

Remember how to integrate an exponential keynote? :)\[\Large\rm \int\limits e^{at}dt=\frac{1}{a}e^{at}\]We just divide by that coefficient on the t.

zepdrix (zepdrix):

Keynooooote, where you at? Play us some music! Jam those notes!

zepdrix (zepdrix):

t.t

OpenStudy (anonymous):

I am right here. I went out . Am just confused

zepdrix (zepdrix):

About which part? :U

zepdrix (zepdrix):

I mean, there is still a bit of work to do in order to wrap this up. But I want to know what you're confused on up to this point.

zepdrix (zepdrix):

Oh oh we also want to factor a negative out of each term in the exponential so it looks like this:\[\Large\rm \mathscr{L}\left(4^t\right)=\int\limits\limits_0^{\infty} e^{-(s-\ln4)t}dt\]

OpenStudy (anonymous):

Gimme a minute. Lemme try to follow along one more time

OpenStudy (anonymous):

How do you do the integration

zepdrix (zepdrix):

You really want to be comfortable with this idea:\[\Large\rm \int\limits\limits e^{\color{#CC0033}{a}t}dt=\frac{1}{\color{#CC0033}{a}}e^{\color{#CC0033}{a}t}\]Yes, you can do a u-substitution if you prefer, but that's really burdensome. You gotta get quick with these :)

zepdrix (zepdrix):

Notice that we got our problem into this form:\[\Large\rm \mathscr{L}\left(e^{at}\right)=\int\limits\limits_0^{\infty} e^{\color{#CC0033}{-(s-a)}t}dt\]So using that rule for integrating exponential we get:\[\Large\rm =\frac{1}{\color{#CC0033}{-(s-a)}}e^{\color{#CC0033}{-(s-a)}t}\quad|_0^{\infty}\]

zepdrix (zepdrix):

So for our problem:\[\Large\rm \int\limits\limits\limits_0^{\infty} e^{\color{#CC0033}{-(s-\ln4)}t}dt\quad=\quad \frac{1}{\color{#CC0033}{-(s-\ln4)}}e^{\color{#CC0033}{-(s-\ln4)}t}\quad|_0^{\infty}\]

zepdrix (zepdrix):

From there we have to evaluate it at the limits.

OpenStudy (anonymous):

Okay that's better coz I was getting something different